You have a bag of marbles. 5 yellow, 5, green, 5



purple, 5 black, and 5 red. What is the



probability of pulling out a purple marble,



replacing it, then pulling out a black marble?



5



6



7

1
Expert's answer
2022-04-22T11:20:01-0400

"P=\frac{5}{25}\frac{5}{25}=\frac{1}{25}=0.04."